An open train car, with a mass of 2030 kg, coasts along a horizontal track at the speed 2.33 m/s. The car passes under a loading chute and, as it does so, gravel falls vertically into it for 3.41 s at the rate of 425 kg/s. What is the car\'s speed after the loading is completed? Ignore rolling friction (as unrealistic as that may be).
Answer #1

momentum conservation

Pinitial = Pfinal

weight gained by car = 425*3.41= 1449.25 kg

2030*2.33 = (2033+425*3.41)*V

=> V = 1.35 m/s